# MagicClay: Sculpting Meshes With Generative Neural Fields

### v1

@@ Abstract.
The recent developments in neural fields have brought phenomenal capabilities to the field of shape generation, but they lack crucial properties, such as incremental control — a fundamental requirement for artistic work. Triangular meshes, on the other hand, are the representation of choice for most geometry related tasks, offering efficiency and intuitive control, but do not lend themselves to neural optimization. To support downstream tasks, previous art typically proposes a two-step approach, where first a shape is generated using neural fields, and then a mesh is extracted for further processing. Instead, in this paper we introduce a hybrid approach that maintains both a mesh and a Signed Distance Field (SDF) representations consistently. Using this representation, we introduce MagicClay — an artist friendly tool for sculpting regions of a mesh according to textual prompts while keeping other regions untouched. Our framework carefully and efficiently balances consistency between the representations and regularizations in every step of the shape optimization; Relying on the mesh representation, we show how to render the SDF at higher resolutions and faster. In addition, we employ recent work in differentiable mesh reconstruction to adaptively allocate triangles in the mesh where required, as indicated by the SDF. Using an implemented prototype, we demonstrate superior generated geometry compared to the state-of-the-art, and novel consistent control, allowing sequential prompt-based edits to the same mesh for the first time.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
Figure 1. MagicClay introduces a novel sculpting tool, employing a hybrid mesh-SDF representation. The user selects a region (middle) of an input mesh (left), inputs a text-prompt (bottom), and the region automatically grows to match the prompt (right), while the rest of the shape is unchanged, and the mesh remains topologically valid. This enables operations such as sequential semantic mesh editing.
@@ 1. Introduction
The field of 3D shape generation has always been heavily dependent on the representations it uses for the shapes. Recent Neural Field-based representations (i.e., NeRFs [Mildenhall et al., 2021] or SDFs [Park et al., 2019; Chen and Zhang, 2019]), have shown remarkable progress to the task [Poole et al., 2022; Wang et al., 2023a] in a very short time. These representations are robust to noisy losses, and are naturally well-suited for neural frameworks, yielding impressive results and avoiding local minima. On the other hand, these representations are expensive to evaluate (limited by volumetric rendering resolutions), and lack acutely in control (such localized edits or even shape smoothing).
@@
In contrast, triangular meshes are the dominant representation for most 3D applications in industry. This is because meshes are inexpensive, consistent, and intuitive. For instance, when evolving a mesh-based shape, an artist performs edits on geometry, textures, or even topology, and expects any additional updates to retain these attributes. This is not possible currently using Neural Field-based representations. Unfortunately, while the adaptive, or non-uniform, nature of meshes is perhaps their greatest advantage, it is also the reason they are not preferred by current generative frameworks. The sparse gradients induced by meshes tend to limit the ability of optimizations to achieve large deformations in a stable manner.
@@
For this reason, many works turn to implicit functions for coarse generation as a first step, and to meshes in a second step, for the purpose of finer details, or downstream editability. However, as we demonstrate, two-stage pipelines are prone to local minima, and additionally cannot be extended to edit an existing mesh with pre-computed UVs for example.
@@
In this paper, we present MagicClay - a shape evolution and editing framework, based on a hybrid implicit-explicit representation, benefiting from the best of both worlds. MagicClay optimizes a mesh and an SDF jointly in every step throughout the generation process, and introduces a novel sculpting tool for 3D generative modeling. Sculpting is a common approach for 3D modeling, common to commercial 3D modeling software [Blender, 2024; ZBrush, 2024; SubstanceModeler, 2024]. While sculpting currently requires a lot of time and expertise, our new tool allows artists to select a region on a mesh to be modified, provide a textual prompt, and hallucinate an updated region (See Figure 1). As we demonstrate, in addition to the contribution to control, our hybrid representation also benefits computational efforts and overall geometric quality, as various priors can be placed more intuitively on the two representations.
@@
The key technical challenge of the hybrid approach is keeping the two representations synced efficiently. To achieve this, we differentiably render both representations from various angles, and require consistency in RGB renders, opacity and normal maps. Furthermore, we rely on the in-sync-mesh representation to render the SDF at higher resolutions and faster; Instead of the hundreds of samples per rays, we localize the SDF sampling around the mesh surface, and use as little as three samples. Critically, evolving a mesh consistently and stably is an additional challenge. In terms of resolution, a coarse mesh would not be expressive enough for novel details, and a fine mesh is expensive and unstable. Hence, an adaptive tessellation is required, that evolves along with the shape where required. We rely on recent developments in differentiable mesh reconstruction [Barda et al., 2023] to achieve dynamic mesh topology updates, including face splitting, edge collapse, and edge flips. To texture the mesh despite changes in its topology, we contribute a new strategy based on triangle supersampling. Importantly, using this layer, we can maintain mesh properties throughout the optimization.
@@
As we demonstrate, our hybrid approach allows localized and sequential mesh editing operations, preserving mesh topology and information on one hand, while allowing radical and semantic evolution on the other. In addition, we show overall higher generated geometric quality, thanks to the priors the two representations impose on each other. The hybrid approach and sculpting tool demonstrate how the merits of both leading representations can be combined. In essence, our framework brings the recent and future breakthroughs in neural shape generation closer to artistic workflows, where work is in incremental steps, giving the artist precision and control over the end result, but with unprecedented expressiveness.
@@
Figure 2. Overview of the hybrid optimization. We jointly optimize a mesh, an SDF and a shared appearance MLP according to an input prompt. We can either optimize the full geometry, or only a user-selected portion of the mesh for an iterative 3D modeling workflow. We start by differentiably rendering both representations, and enforce their consistency. As they are kept in sync, we use the mesh to efficiently sample volumetric rays to render hi-res maps from the SDF in a memory-efficient manner. In addition to the standard SDS loss, the high-res renderings allows using high quality VSD losses [Wang et al., 2023a] to evolve the SDF. We keep the mesh surface and the SDF in sync via multi-view consistency constraints on the RGB pixels, the image opacity and the surface normals. The mesh local topology is updated according to the SDF using ROAR [Barda et al., 2023], splitting triangles where geometry is created, and collapsing edges where needed. Additionally, we leverage representation-specific losses to regularize the optimization: an Eikonal loss on the SDF and a smoothness loss on the mesh.
@@ 2. Related Work
In their seminal work, DreamFusion, pool et al. [Poole et al., 2022] show that Text-to-Image diffusion models can be used to provide gradients to optimize a Neural Radiance Field (NeRF) via Score Distillation Sampling (SDS). Magic3D [Lin et al., 2023] achieves better quality by using a two stage approach: the first stage is similar to DreamFusion, and they note that the quality of the generated object is limited by the high cost of performing volumetric rendering for high resolutions images. The second stage uses a differentiable mesh representation to further refine the generated object, as differentiably rendering meshes in high resolution is significantly cheaper in both time and memory. Magic123 [Qian et al., 2023] further improves upon Magic3d by using both 3d and 2d diffusion priors. ProlificDreamer [Wang et al., 2023a] proposes an improvement over SDS, the VSD loss, to drive 3D generation from 2d diffusion priors. Fantasia3d [Chen et al., 2023] and TextMesh [Tsalicoglou et al., 2023] decouple the appearance from the geometry by replacing the NeRF with an SDF, and optimizing a color network separately. TextDeformer [Gao et al., 2023] uses CLIP as prior together with a novel gradient smoothing technique based on mesh Jacobians to deform meshes according to a text prompt.
@@
The choice of using two stages in Magic3D [Lin et al., 2023] highlights the tradeoffs involved in choosing the right 3D representation for 3D generative models. While implicit functions are well suited for coarse generation because they allow topology updates, meshes can be rasterized very efficiently at a high resolution to get fine details in a second step. However, two-stage pipelines are prone to local minima and crucially, cannot be extended to edit an existing mesh with pre-computed UVs. In contrast, we jointly optimise a hybrid SDF and mesh representation, that can be initialized from an existing mesh and maintain all of its properties during optimization, benefiting from the best of both worlds in a 1-step pipeline : topology updates from the SDF part and fine details from the mesh part.
Focus on local editing: Vox-E [Sella et al., 2023] edit a voxel grid via SDS and use the attention layers to encourage localized edits. However, they can not guarantee where the edit will happen because the localization mechanism is based on soft attention.
Concurrent work: DreamCraft3D [Sun et al., 2023] builds on the recent SDS works by fine-tuning the diffusion model during the generative process using DreamBooth [Ruiz et al., 2022]. Of note, Instant3D [Li et al., 2023] generates a 3D shape in a single forward pass, without require any costly optimization. While it does not allow for local artistic controls similar to the sculpting application we present, we are excited to leverage ideas from this research direction to accelerate our method in the future.
Hybrid Representations.
@@
There is no ubiquitous representation in 3D, as there exist in 2D for images, thus several representations exist and have been combined for diverse 3D tasks. The plethora of representations and combination show that there is no one-fit-for-all solution. In this work, we introduce a hybrid representation specialized for generative modeling and focus the related work on hybrids most relevant to this paper. Poursaeed et al. [Poursaeed et al., 2020] uses a coupling of implicit and explicit surface representation for generative 3D modeling, kept in sync by 3D losses. NerfMeshing [Rakotosaona et al., 2023] proposes a improved meshing pipeline for NeRFs. In contrast to  [Poursaeed et al., 2020], the coupling is not achieved via coupled regularization losses, but explicitely enforced by projection layers from the SDF to the mesh. Finally, DmTeT [Shen et al., 2021] proposes deep marching Tetrahedra as a hybrid representation for high-resolution 3D Shape synthesis, notably used in the concurrent work Magic3D [Lin et al., 2023]. Our method uses both a set of regularization losses, as well as a dynamic projection layer based on ROAR [Barda et al., 2023] to keep the SDF and mesh part in sync.
Traditional approaches for sculpting meshes
@@
Many commercial tools employ digital sculpting metaphor for 3D modeling, such as Zbrush [ZBrush, 2024], Mudbox [Autodesk, 2024], or SubstanceModeler [SubstanceModeler, 2024]. Motivated by these workflows, geometry processing research focused on improving interactive techniques such as mesh deformation [Jacobson et al., 2014], mesh blending for cut-and-paste [Biermann et al., 2002], local parameterization for adding surface details [Schmidt et al., 2006], symmetry-guided autocompletion [Peng et al., 2018], and version control for collaborative editing [Salvati et al., 2015]. Despite these advances, 3D modeling remains to be only accessible to experts. As an alternative, example-based approaches have been proposed to democratize 3D modeling tools by using existing geometry from a database of stock 3D models to assemble new shapes from parts [Funkhouser et al., 2004]. Subsequent methods have built statistical models over part assemblies [Kalogerakis et al., 2012], and allow high-level semantic control for deformations [Yumer et al., 2015]. Despite their accessibility, these tools are often restricted in their domain, and often rely on heavy annotation of stock assets, and thus have received limited use by professional modelers. In this paper, we utilize pretrained 2D generative data prior to enable semantic controls for local and iterative modeling workflow without the need of preannotated 3D stock data.
@@ 3. Method
Given a mesh, a user-highlighted surface region, and a text prompt that describes the desired target, MagicClay optimizes the shape of the selected region so that the resulting mesh matches the target. To drive the shape optimization, we follow current literature and use the Score Distillation Sampling (SDS) technique [Poole et al., 2022] with differentiable rendering to leverage on text-conditioned 2D diffusion and guide the shape optimization. This approach, however, does not perform well when operated on meshes. Meshes are driven by sparse and irregular samples (vertices), and their connectivity mandates a stable and smooth deformation, avoiding self-intersections and flip-overs. For this reason, we employ a neural Signed Distance Field (SDF) to drive the mesh shape optimization and topology updates. We thus propose a hybrid representation that captures both a Signed Distance Field (SDF) and the surface, gaining from the advantages of both worlds. While the SDF allows guiding the shape towards larger-scale complex changes, the mesh allows to capture fine details, and localize control to the user-highlighted surface region.
In this section we provide details on the hybrid SDF/Mesh representations (Sec. 3.1), how it can be efficiently optimized with SDS guidance (Sec. 3.2), how to effectively use surface and volumetric priors (Sec. 3.3), and how to update the mesh topology during optimization (Sec. 3.4). Figure 2 overviews the full pipeline.

### v2

@@ Abstract.
The recent developments in neural fields have brought phenomenal capabilities to the field of shape generation, but they lack crucial properties, such as incremental control — a fundamental requirement for artistic work. Triangular meshes, on the other hand, are the representation of choice for most geometry-related tasks, offering efficiency and intuitive control, but do not lend themselves to neural optimization. To support downstream tasks, previous art typically proposes a two-step approach, where first, a shape is generated using neural fields, and then a mesh is extracted for further processing. Instead, in this paper, we introduce a hybrid approach that maintains both a mesh and a Signed Distance Field (SDF) representations consistently. Using this representation, we introduce MagicClay — a tool for sculpting regions of a mesh according to textual prompts while keeping other regions untouched. Our method is designed to be compatible with existing mesh sculpting workflows. The user sculpts the desired shape using the existing brushes and our pipeline then evolves the geometry and triangulation of the selected mesh part according to the given textual prompt. This process operates on the original mesh while preserving its meta-data Our framework carefully and efficiently balances consistency between the representations and regularizations in every step of the shape optimization. Relying on the mesh representation, we show how to render the SDF at higher resolutions and faster. In addition, we employ recent work in differentiable mesh reconstruction to adaptively allocate triangles in the mesh where required, as indicated by the SDF. Using an implemented prototype, we demonstrate superior generated geometry compared to the state-of-the-art and novel consistent control, allowing sequential prompt-based edits to the same mesh for the first time. We will release the code upon acceptance.
submissionid: 616
Figure 1. Top: the user selects a region of an input mesh, coarsely sketches the intended edit, inputs a text-prompt, and MagicClay grows the region automatically to match the prompt, while the rest of the shape and existing textures remain unchanged, and the mesh remains topologically valid (no non-manifold edges or vertices). Importantly, due to working directly on the input mesh, these edits can be made sequentially. Bottom: MagicClay preserves attributes from the input mesh such as texture, rig, and tessellation, e.g. allowing to transfer an animation from source to edited shape.
@@ 1. Introduction
The field of 3D shape generation has always been heavily dependent on the representations it uses for the shapes. Recent neural field-based representations (i.e., NeRFs [Mildenhall et al., 2020] or SDFs [Park et al., 2019; Chen and Zhang, 2019]), have shown remarkable progress to the task [Poole et al., 2023; Wang et al., 2023a] in a very short time. These representations are robust to noisy losses and are naturally well-suited for neural frameworks, yielding impressive results and avoiding local minima. On the other hand, these representations are expensive to evaluate (limited by volumetric rendering resolutions) and lack acutely in geometric control, such as allowing users to localize their edits or leverage surface-based (e.g., smoothness) priors. For instance, when iterating on a design of a mesh-based 3D model, as artists alter geometry, textures, or topology, they expect any additional updates to retain previously assigned attributes. This level of control is currently not feasible with the neural field-based representations.
@@
In contrast, triangular meshes provide such control and are indeed the current dominant representation for most 3D applications in the industry, as they are computationally inexpensive, consistent, and intuitive. On the other hand, while the adaptive, non-uniform nature of meshes is one of their greatest advantages, it is also the reason they are not widely used in current generative frameworks. The sparse gradients induced by meshes tend to limit the ability of optimizations to achieve large deformations in a stable manner.
@@
For this reason, many works [Lin et al., 2023] turn to a two-step process, where first implicit functions are used for coarse generation, and then are converted to meshes in a second step for the purpose of finer details or downstream editability. However, as we demonstrate, two-stage pipelines are prone to local minima and cannot be extended to the task of editing an existing mesh with pre-existing UVs and textures, for example.
@@
In this paper, we present MagicClay - a shape editing framework based on a hybrid implicit-explicit representation. MagicClay optimizes an input mesh (possibly textured) and its SDF jointly at every step of the generation process and leverages consistency and representation-specific priors, benefiting from the best of both worlds. This approach leads to higher quality generation output and also enables a higher level of control, allowing the users to sequentially sculpt local generative details in regions marked as editable, guided by the user-provided prompt. Sculpting is a common approach used in 3D modeling software [Blender, 2024; ZBrush, 2024; SubstanceModeler, 2024]. While sculpting currently requires a lot of time and expertise, our novel tool offers unprecedented control by allowing artists to select a region on a mesh to be modified, provide a textual prompt, and hallucinate an updated region (Figure 1).
@@
The key technical challenge of the hybrid approach is keeping the two representations synced efficiently. To achieve this, we differentiably render both representations from various angles and require consistency in RGB renders, opacity, and normal maps. Furthermore, we rely on the in-sync-mesh representation to render the SDF at higher resolutions and faster; instead of the hundreds of samples per ray required in most previous techniques [Poole et al., 2023], we localize the SDF sampling around the mesh surface and use as little as three samples. Critically, optimizing a mesh consistently and stably is an additional challenge. In terms of resolution, a coarse mesh would not be expressive enough for novel details, and a fine mesh is expensive and unstable. Hence, an adaptive tessellation is required, that changes along with the shape where needed. We rely on recent developments in differentiable mesh reconstruction [Barda et al., 2023] to achieve dynamic mesh topology updates, including face splitting, edge collapse, and edge flips. To texture the mesh despite changes in its topology, we contribute a new strategy based on triangle supersampling. Importantly, using this layer, we can maintain mesh properties (e.g, vertex groups for animation) throughout the optimization, and through local mesh topology changes.
@@
As we demonstrate, our hybrid approach allows localized and sequential mesh editing operations, allowing the user to preserve existing mesh triangulation and information in some regions, while allowing radical and semantic changes in other regions. In addition, we show overall higher generated geometric quality as compared to using an implicit representation alone, thanks to the priors the two representations impose on each other. By combining the merits of implicit and explicit representations to enable a novel generative sculpting tool, we bring the neural shape generation pipeline closer to the artistic workflows, allowing for incremental editing steps and providing the artist with precision and control over the end result.
To summarize, our main contributions are :
  • A new hybrid representation that brings the benefits of SDFs and meshes together : SDFs are more robust to noisy gradients (see Figure 3) and meshes allow for surface-based losses and localization of edits. Both representations are kept consistent via multi-view consistency losses and adaptive remeshing. In Figure 4 and Table 1, we demonstrate in the task of unconditional text-to-3D that MagicClay produces higher-quality meshes than other representations using similar generative techniques.
  • An application of the new representation to localized mesh editing, which brings the prowess of generative techniques to mesh sculpting, while critically preserving the part of the geometry that is not selected for editing, including its texture, tessellation, and rigging parameters. We show in Figure 8 that our method produces higher-quality and more localized edits than competing methods.
@@
Figure 2. Overview of the hybrid optimization. We jointly optimize a mesh, an SDF and a shared appearance MLP according to an input prompt. We can either optimize the full geometry, or only a user-selected portion of the mesh for an iterative 3D modeling workflow. We can also preserve existing textures on non-selected part of the mesh, or have the diffusion model generate textures for the full mesh. We start by differentiably rendering both representations, and enforcing their consistency. As they are kept in sync, we use the mesh to efficiently sample volumetric rays to render hi-res maps from the SDF in a memory-efficient manner. Applying SDS-type losses on these Hi-res renderings allows for capturing finer details. We sync the mesh and the SDF via multi-view consistency constraints on the RGB pixels, the image opacity, and the surface normals. The mesh local topology is updated according to the SDF using ROAR [Barda et al., 2023], splitting triangles where geometry is created and collapsing edges where needed. Additionally, we leverage representation-specific losses to regularize the optimization: an Eikonal loss on the SDF and a smoothness loss on the mesh.
@@ 2. Related Work
In their seminal work, DreamFusion, [Poole et al., 2023] show that Text-to-Image diffusion models can be used to provide gradients to optimize a Neural Radiance Field (NeRF) via Score Distillation Sampling (SDS). Magic3D [Lin et al., 2023] achieves better quality by using a two-stage approach: the first stage is similar to DreamFusion, and they note that the quality of the generated object is limited by the high cost of performing volumetric rendering for high-resolution images. The second stage uses a differentiable mesh representation to refine the generated object further, as differentiably rendering meshes in high resolution is significantly cheaper in both time and memory. Magic123 [Qian et al., 2024] further improves upon Magic3D by using both 3D and 2D diffusion priors. ProlificDreamer [Wang et al., 2023a] proposes an improvement over SDS, the VSD loss, to drive 3D generation from 2d diffusion priors. Fantasia3D [Chen et al., 2023] and TextMesh [Tsalicoglou et al., 2024] decouple the appearance from the geometry by replacing the NeRF with an SDF, and optimizing a color network separately. TextDeformer [Gao et al., 2023] uses CLIP as prior together with a novel gradient smoothing technique based on mesh Jacobians to deform meshes according to a text prompt.
@@
The choice of using two stages in Magic3D [Lin et al., 2023] highlights the tradeoffs involved in choosing the right representation for 3D generative models. While implicit functions are well suited for coarse generation because they allow topology updates, meshes can be rasterized very efficiently at a high resolution to get fine details in a second step. However, two-stage pipelines are prone to local minima and crucially, cannot be extended to edit an existing mesh with pre-computed UVs. In contrast, we jointly optimize a hybrid SDF and mesh representation, that can be initialized from an existing mesh and maintain all of its properties during optimization, benefiting from the best of both worlds in a 1-step pipeline : topology updates from the SDF part and fine details from the mesh part.
Finally, Instant3D [Li et al., 2024] generates a 3D shape in a single forward pass without requiring any costly optimization. Though faster and higher-quality, Instant3D does not allow sculpting an existing mesh like MagicClay.
Generative local editing.
A recent line of work deals with locally editing a 3D scene. Unlike 2D images, selecting a region in 3D to be edited is not straightforward for implicit functions. Most of these works solely rely on soft attention between the text and the renderings to semantically understand the area to be edited, and they vary with their choice of 3D representation. Vox-E [Sella et al., 2023] and DreamEditor [Zhuang et al., 2023] encode an SDF respectively with a voxel grid and an MLP. Instruct-NerftoNerf [Haque et al., 2023], LatentNerf  [Metzer et al., 2023], SKED [Mikaeili et al., 2023], MVEdit [Chen et al., 2024] use a NeRF. In these methods, the user selects the region to edit via the prompt, except for SKED [Mikaeili et al., 2023], which additionally utilizes guiding sketches from different views. FocalDreamer [Li et al., 2023] and Progressive3D [Cheng et al., 2024] use an implicit representation with SDS and additionally allow selection of 3D region(s) with losses encouraging localized changes. The output mesh is reconstructed from the edited SDF using DMTeT, thus losing all mesh properties in the non-editable regions. In contrast, our method exposes a mesh to the user, on which it is trivial to select a region and guarantees that the unselected part will be unaltered.
@@
Hybrid Representations. unlike for 2D images, there is no ubiquitous representation in 3D, , and several representations exist and have been combined for diverse 3D tasks, suggesting that there is no one-fit-for-all solution. In this work, we introduce a hybrid representation specialized for generative modeling and focus the related work on hybrids most relevant to this paper.  [Poursaeed et al., 2020] uses a coupling of implicit and explicit surface representations for generative 3D modeling, kept in sync by 3D losses. NerfMeshing [Rakotosaona et al., 2024] proposes an improved meshing pipeline for NeRFs. Finally, DmTeT [Shen et al., 2021] proposes deep marching Tetrahedra as a hybrid representation for high-resolution 3D Shape synthesis, notably used in the concurrent work Magic3D [Lin et al., 2023]. Our method uses both a set of regularization losses, as well as a dynamic projection layer based on ROAR [Barda et al., 2023] to keep the SDF and mesh part in sync.
@@
Traditional approaches for sculpting meshes Many commercial tools employ the digital sculpting metaphor for 3D modeling, such as Zbrush [ZBrush, 2024], Mudbox [Autodesk, 2024], or SubstanceModeler [SubstanceModeler, 2024]. Motivated by these workflows, geometry processing research has focused on improving interactive techniques such as mesh deformation [Jacobson et al., 2014], mesh blending for cut-and-paste [Biermann et al., 2002], local parameterization for adding surface details [Schmidt et al., 2006], symmetry-guided autocompletion [Peng et al., 2018], and version control for collaborative editing [Salvati et al., 2015]. Despite these advances, 3D modeling remains only accessible to experts. As an alternative, example-based approaches propose to democratize 3D modeling tools by using existing geometry from a database of stock 3D models to assemble new shapes from parts [Funkhouser et al., 2004]. Subsequent methods have built statistical models over part assemblies [Kalogerakis et al., 2012], and allow semantic control for deformations [Yumer et al., 2015]. Despite their accessibility, these tools are often restricted in their domain, and rely on heavy annotation of 3D assets, and thus have received limited use by professional modelers. In this paper, we utilize pre-trained 2D generative data priors to enable semantic controls for local and iterative modeling workflow without the need for pre-annotated 3D data.
@@ 3. Method
Given a mesh, a user-highlighted surface region, and a text prompt that describes the desired target, MagicClay optimizes the shape of the selected region so that the resulting mesh matches the target. To drive the shape optimization, we follow current literature and use the Score Distillation Sampling (SDS) technique [Poole et al., 2023] with differentiable rendering to leverage on text-conditioned 2D diffusion and guide the shape optimization. This approach, however, does not perform well when operated on meshes. Meshes are driven by sparse and irregular samples (vertices), and their connectivity mandates a stable and smooth deformation, avoiding self-intersections and flip-overs. For this reason, we employ a neural Signed Distance Field (SDF) to drive the mesh shape optimization and topology updates. We thus propose a hybrid representation that captures both a Signed Distance Field (SDF) and the surface, gaining from the advantages of both worlds. While the SDF allows guiding the shape toward larger-scale complex changes, the mesh allows localized control of the user-highlighted surface region.
In this section, we provide details on the hybrid SDF/Mesh representations (Sec. 3.1), how it can be efficiently optimized with SDS guidance (Sec. 3.2), how to effectively use surface and volumetric priors (Sec. 3.3), and how to update the mesh topology during optimization (Sec. 3.4). Figure 2 overviews the full pipeline.